At an annual rate of return of 8%, what is the future worth of $100 at the end of four years

a. $130

b. $132

c. $135

d. $136

e. $140
Answer #1

Answer : The answer is option d.

Rate of interest (r) = 8% = 0.08

Principal (P) = $100

Time (t) = 4 years

n = Per year interest rate compounding time.

Future worth = A

Now,

A = P (1 + r/n)^nt

=> A = 100 (1 + 0.08/1)^(1 * 4)

=> A = 100 * 1.36

=> A = $136

Therefore, option d is correct.